OpenStudy (anonymous):

Determine the period, to the nearest tenth of radian, for the following function: y = 4 cos(4x + 6) - 1

OpenStudy (mertsj):

\[y=4\cos (4x+6)-1=4\cos 4(x+\frac{3}{2})-1\]

OpenStudy (mertsj):

If you want to find the period, you should use this equation: \[coefficient of x = \frac{2\pi}{period}\] \[4=\frac{2\pi}{p}\]

OpenStudy (mertsj):

\[4p=2\pi\] \[p=\frac{\pi}{2}\]

OpenStudy (anonymous):

ok can u help me with this too If the maximum value of y = a cos x + 2.4 is 12.6, then the value of a, to the nearest tenth, is

OpenStudy (mertsj):

Typically a cosine graph would go up to 1 and down to -1. Since this one is raised 2.4, it's centerline is y = 2.4 instead of y=0. It goes up to 12.6 which is 10.2 units above its centerline so its amplitude must be 10.2